# Operator Learning with Domain Decomposition for Geometry Generalization in PDE Solving

### v1

@@ 3.3 Schwarz Neural Inference
Normalization. During inference on an arbitrary decomposed subdomain, the range of geometry and boundary conditions may differ from that of the generated training data. We thus leverage the symmetry properties of PDEs to handle this mismatch. More specifically, we can directly apply transformations $T:\mathcal{P}\times\mathcal{H}\to\mathcal{P}\times\mathcal{H}$ such as spatial translation and scaling laws to transform a local problem outside our training range - geometry or function values - into the training range. Note that the transformations have to be extended to any coefficient fields or source term if they are also involved in the symmetry. After neural operator inference, the resulting solution function will be transformed back by a proper inverse transformation $\tilde{T}:\mathcal{U}\to\mathcal{U}$. We implement these transformations as preprocessing and postprocessing steps in the inference pipeline.
Time Complexity. Suppose the single inference time of local operator and the number of iterations are denoted as $b$ and $N$. Let $v,e,K$ denote the number of vertices, edges and subdomains respectively. Our Algorithm 1 consists of two main parts: mesh partition using the METIS algorithm, the time complexity of which is approximately $O(v+e+K\log K)$ Karypis and Kumar, (1997); iterative scheme using the additive Schwarz method with a time complexity roughly $O(bKN)$. Therefore, the overall time complexity of our algorithm can be approximated as $O(v+e+K\log K+bKN)$ [^1] With neural operators implementing linear transformers, e.g., GNOT applied in this work, $b=O(\frac{v}{K}).$. While providing an exact time complexity analysis for FEM can be challenging due to the complexity and variability of different problem setups, it is worth noting that FEM is generally considered to be computationally demanding.
@@ 3.4 Theoretical Results
Let $u$ and $\tilde{u}$ denote the solution obtained by classical additive Schwarz method given Eq. (2) and SNI in Algorithm 1, respectively, with the same initial condition $u^{0}=\tilde{u}^{0}$. Assume $\parallel\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}(\cdot)-\mathcal{S}_{k}(\cdot)\parallel<c$ for all $k$, and if the classical algorithm converges, then we have:
  • Convergence: SNI converges to a fixed point;
  • Error bound: there exists a constant $c^{\prime}$ (depending on $c$) such that $\parallel\tilde{u}^{n}-u^{n}\parallel<c^{\prime}$.
The theorem suggests that if our learned local operator maintains a uniform error bound, the algorithm converges and exhibits a minimal approximation error. See Appendix A.3 for a proof.
@@ 4.1 Experimental Setup
Datasets. To demonstrate the scalability and superiority of our method, we construct several datasets on multiple PDEs. We also extend our framework to a time-dependent problem, heat conduction. To aggregate training sets, we generate random simple polygons bounded by the unit square $[-0.5,0.5]^{2}\subset\mathbb{R}^{2}$. Boundary/initial conditions and coefficient functions are piecewise linear functions determined by random values within $[0,1]$. For each of the following problems, we test on datasets based on three different domains A, B and C shown in Figure 2. Details of these datasets are given in Appendix A.7.1.
@@
Baseline. We compare our method with a direct inference of the trained neural operator on domains shifted and scaled to $[-0.5,0.5]^{2}$.
Evaluation Protocol. The evaluation metric we utilize is the mean $l_{2}$ relative error. See Appendix A.7.2 for details.
@@ 4.2 Main Results and Analysis
Equation Domain GNOT SNI
Laplace2d-Dirichlet A (21.53$\pm$2.46)e-2 (2.15$\pm$0.58)e-2
B (22.39$\pm$1.89)e-2 (2.10$\pm$0.40)e-2
C (28.45$\pm$2.82)e-2 (2.10$\pm$0.86)e-2
Laplace2d-Mixed A (10.74$\pm$0.79)e-2 (6.36$\pm$3.70)e-2
B (10.74$\pm$0.79)e-2 (7.00$\pm$1.32)e-2
C (37.53$\pm$5.71)e-2 (6.22$\pm$1.28)e-2
Darcy2d A (15.65$\pm$1.25)e-2 (8.60$\pm$1.61)e-2
B (63.42$\pm$3.01)e-2 (7.94$\pm$1.65)e-2
C (166.88$\pm$7.74)e-2 (5.35$\pm$0.64)e-2
Heat2d A (11.5$\pm$0.60)e-2 (5.28$\pm$0.16)e-2
B (31.1$\pm$14.8)e-2 (11.5$\pm$1.70)e-2
C (22.3$\pm$14.6)e-2 (5.76$\pm$0.27)e-2
@@
The main results for all datasets are shown in Table 1. More details and hyperparameters are summarized in Appendix A.7.2 due to limited space. Based on these results, we have the following observations.
@@
(d) Darcy2d domain A
(e) Darcy2d domain B
(f) Darcy2d domain C
Figure 3: Comparison between the $l_{2}$ relative errors from SNI (blue), GNOT direct inference (orange) and validation (red) on two different PDEs (Laplace2d-Dirichlet and Darcy2d) upon three domains (A, B and C) with different numbers of training samples. The results of SNI and GNOT direct inference are presented based on 100 inferences with different boundary conditions. The best validation errors during training are also provided as a reference.
@@
Time-dependent Problems. There is a natural way to extend our framework to time-dependent problems Li and Cai, (2015) where a space-time decomposition is constructed by taking the product of a spatial decomposition and a temporal decomposition. We train a neural operator that can predict heat conduction on multiple time steps and the same SNI is applied during inference on this 3d problem. Our framework works well on this problem and reduce prediction error by 54.1%-74.2%. This demonstrates the potential of our framework to handle time-dependent problems. We refer to Appendix A.4 for detailed implementation.
Data Efficiency. The exploration results on data efficiency of SNI are shown in Figure 3, implying the following observations: (1) At all abundances of data, the $l_{2}$ relative errors of SNI are significantly lower than those of GNOT direct inferences; (2) Errors of SNI are comparable to or even lower than validation errors at large data volumes. (3) SNI requires much smaller datasets to achieve comparable results to GNOT direct inference. Overall, these results demonstrate that SNI has substantial advantages in terms of data efficiency. Our proposed framework possesses remarkable ability to extract more insights from limited data and scale more effectively as data volumes increase. More supplementary results are provided in Appendix A.7.3.
@@ 4.3 Ablation Experiments
Validation Domain A Domain B Domain C
No Data Aug 3.79e-2 (4.38$\pm$1.58)e-2 (2.95$\pm$0.58)e-2 (2.75$\pm$0.95)e-2
Rotation Only 2.50e-2 (2.15$\pm$0.58)e-2 (2.10$\pm$0.40)e-2 (2.10$\pm$0.86)e-2
Rotation + Scale [0.2, 1] 5.31e-2 (3.74$\pm$1.20)e-2 (3.37$\pm$0.50)e-2 (3.41$\pm$0.43)e-2
[0.5, 1] 3.62e-2 (2.71$\pm$0.48)e-2 (3.73$\pm$0.61)e-2 (3.21$\pm$0.57)e-2
[0.8, 1] 2.86e-2 (1.87$\pm$0.43)e-2 (3.30$\pm$0.71)e-2 (2.78$\pm$0.82)e-2
@@
Hyperparameter Exploration. The number of partitions ($K$), the depth of extension ($d$) and step size ($\tau$) are the key main hyperparameters that can affect the performance of SNI. Based on the results presented in Figure 4, the factors analyzed have no significant impact on the accuracy of our algorithm, but they do influence the convergence rate. Specifically, increasing the number of partitions leads to a smaller $l_{2}$ relative error but slower convergence. Once the partition number surpasses 20, the algorithm’s final performances become comparable. However, when using 50 partitions, the inference process does not converge in 5000 iterations. Regarding the depth of extension, it does not affect the performance on the tested domains. The convergence curves for depth of extension 1, 2, 4, and 8 are nearly identical. When it comes to $\tau$, a larger value results in faster convergence. However, it is important to note that there exists a maximum limit $1/K$ beyond which $\tau$ cannot be set.
Data Augmentation Exploration. To explore the effects of data augmentation, we compare the performances of models trained with different degrees of data augmentation for Laplace2d-Dirichlet demonstrated in Table 2. For models trained without data augmentation, the variance of performances on domains is large, ranging from $2.75\%$ to $4.38\%$. Specifically, it reports a $4.38\pm 1.58\%$ $l_{2}$ relative error on domain A, while this error can be reduced to $1.87\pm 0.43\%$ with a rotation+[0.8, 1] scaling augmentation. While rotation can generally be beneficial, the effectiveness of scaling can sometimes be limited or even detrimental. Hence, it is important to apply data augmentation with caution and consider its suitability for different types of PDEs.
@@ 5 Related Work
Operator Learning. The idea of operator learning is first introduced in Lu et al., (2019). This work proposes a notable architecture called DeepONet, which employs a branch network for processing input functions and a trunk network for handling query points. Adopting the trunk-branch architecture and utilizing the attention mechanism,  Hao et al., (2023) develops GNOT to handle irregular mesh, multiple input functions, and different input data types. The high accuracy and versatility makes GNOT the benchmark in our work. In the other direction, Fourier neural operator (FNO) Li et al., 2020b leverages the Fast Fourier Transform (FFT) to learn operators in the spectral domain, and achieves a favorable trade-off between cost and accuracy. Variants of FNO are proposed to reduce computational cost (FFNO in Tran et al., (2021)), handle irregular mesh (GeoFNO in Li et al., (2023)), and improve expressivity (UFNO in Wen et al., (2022)).
@@
Domain Decomposition Methods Applied in Deep Learning. In general, the integration of deep learning and DDMs can be categorized into two groups Heinlein et al., (2021). The first category involves using deep learning techniques to improve the convergence properties or computational efficiency of DDMs. For instance, several methods Heinlein et al., (2020, 2019) have been proposed to reduce the computational cost in adaptive FETI-DP solvers by incorporating deep neural networks (DNNs) while ensuring the robustness and convergence behavior. The second category is centered around the substitution of subdomain solvers in DDMs with neural networks. There have been multiple endeavors to employ PINNs or Deep Ritz methods as alternatives to subdomain solvers or discretization techniques in traditional DDMs Li et al., 2020a ; Li et al., (2019). These approaches leverage the universal approximation capabilities of neural networks to represent solutions of PDEs, subject to specific assumptions regarding the activation function and other factors.
@@
Data Augmentation Techniques in Operator Learning. Different types of data augmentations are proposed to improve the generalization capabilities in operator learning. A Lie point symmetry framework is introduced in Brandstetter et al., (2022), which quantitatively derives a comprehensive set of data transformations, to reduce the sample complexity. Motivated by this approach,  Mialon et al., (2023) learn general-purpose representations of PDEs from heterogeneous data by implementing joint embedding methods for self-supervised learning. An alternative research approach Fanaskov et al., (2023) introduces a computationally efficient augmentation strategy that relies on general covariance and straightforward random coordinate transformations. In general, applying data augmentation techniques for PDE operator learning can be challenging due to the unique nature of PDE theory.
@@ 6 Conclusion and Future Works
Appendix A Appendix / supplemental material
A.1 Background on Domain Decomposition
@@
In this work, we decompose our domain into subdomains and adopt the hybrid formulation of Eq. (1) following (Mathew,, 2008, Section 1.1). A decomposition of $\Omega$ is a collection of open subregions $\{\Omega_{k}\}_{k=1}^{K}$, $\Omega_{k}\subseteq\Omega$ for $k=1,\dots,K$ such that $\bigcup_{k=1}^{K}\overline{\Omega_{k}}=\overline{\Omega}$. This decomposition is referred to as non-overlapping if in addition, $\Omega_{i}\cap\Omega_{j}=\varnothing$ for any $i\neq j$. Alternatively, an overlapping decomposition is one satisfying $\bigcup_{k=1}^{K}\Omega_{k}=\Omega$. Typically, a non-overlapping decomposition is one where subdomains do not intersect with each other in the interior while an overlapping decomposition constructed in practice has overlapping neighboring subdomains.
@@
Given a decomposition of $\Omega$, a hybrid formulation of Eq. (1) is a coupled system of local PDEs on subdomains $\Omega_{k}$ equivalent to Eq. (1) satisfying two requirements. First, the restriction $u_{k}(x)$ of the solution $u(x)$ of Eq. (1) to each domain $\Omega_{k}$ must solve the local PDE, thus ensures that the hybrid formulation is consistent with the original problem in Eq. (1). Second, the hybrid formulation must be well posed as a coupled system of PDEs in the sense of Evans, (2022), i.e. its solution must exist, be unique and depend continuously on given input function and boundary/initial conditions. Intuitively, a hybrid formulation consists of a local problem posed on each subdomain and matching conditions that couples the local problems.
@@
In this work we focus on the earliest and most elementary formulation termed Schwarz hybrid formulation (Mathew,, 2008, 1.2) based on overlapping decomposition and is applicable to a wide class of self-adjoint and coercive elliptic equations. Given an overlapping decomposition, $\partial\Omega_{k}$ can be decomposed into two disjoint parts. One (possibly empty) part $\Gamma_{k}=\partial\Omega_{k}\cap\partial\Omega$ is located in the boundary of $\Omega$ and the global boundary condition should be imposed. The other part $B_{k}=\partial\Omega_{k}\cap\Omega$ is a nonempty artificial boundary from the overlapping decomposition and a Dirichlet boundary condition from the coupling of local problems is imposed.
@@
We refer to Mathew, (2008) for a strict definition. As an illustrative example, assume we have an overlapping decomposition with $K=2$ and consider as the original problem Laplace equation with mixed Dirichlet and Neumann boundary conditions. The following coupled system of two local PDEs is a Schwarz hybrid formulation of the original problem and solving the original equation is equivalent to solving this coupled system.
$\Delta u_{1}&=0\quad\quad\quad\ \text{in}\ \ \Omega_{1}\\ u_{1}&=u_{2}|_{\partial\Omega_{1}}\quad\text{on}\ \ \partial\Omega_{1}\cap\Omega\\ u_{1}&=u_{D}\quad\quad\ \ \text{on}\ \ \partial\Omega_{1}\cap\Gamma_{D}\\ \frac{\partial u_{1}}{\partial n}&=g\quad\quad\quad\ \text{on}\ \ \partial\Omega_{1}\cap\Gamma_{N}\end{aligned}\quad\text{and}\quad\begin{aligned} \Delta u_{2}&=0\quad\quad\quad\ \text{in}\ \ \Omega_{2}\\ u_{2}&=u_{1}|_{\partial\Omega_{2}}\quad\text{on}\ \ \partial\Omega_{2}\cap\Omega\\ u_{2}&=u_{D}\quad\quad\ \ \text{on}\ \ \partial\Omega_{2}\cap\Gamma_{D}\\ \frac{\partial u_{2}}{\partial n}&=g\quad\quad\quad\ \text{on}\ \ \partial\Omega_{2}\cap\Gamma_{N}\\ \end{aligned}$
@@
Based on the Schwarz hybrid formulation, there are various iterative schemes with different parallelism and convergence rate. In the subsequent discussion, our focus is primarily on introducing the additive Schwarz methods (ASM). The ASM is a highly parallel algorithm Mathew, (2008) in solving the coupled system from Schwarz hybrid formulation. We briefly introduce ASM with finite element methods and refer to Gander et al., (2008) and Mathew, (2008) for details.
@@
Assume that under weak formulation of Eq. (1) and finite element space $V$, Eq. (1) has the form $Au=f$ where $A$ is the stiffness matrix. Given an overlapping decomposition $\{\Omega_{i}\}_{k=1}^{K}$ compatible with the finite element space on $\Omega$, we have $V=\sum_{k=1}^{K}V_{k}$ as sum of local finite element subspaces $V_{k}$ on $\Omega_{k}$ and we can define local stiffness matrices $A_{k}:V_{k}\to V_{k}$, restriction operators $\{R_{k}\}_{k=1}^{K}$ restricting $V$ to $V_{k}$ and extension operators $\{R_{k}^{\intercal}\}_{k=1}^{K}$ extending $V_{k}$ to $V$ by zeros extension. We then define operators $P_{k}:V\to V$ by $P_{k}=R_{k}^{\intercal}A_{k}^{-1}R_{k}A$. Additive Schwarz operator is then defined as the sum $P_{\text{ad}}=\sum_{k=1}^{K}P_{k}$. This operator can be show to be self-adjoint and coercive and we have the following equivalence.
@@
We note that the right hand side of Eq. (4) is a preconditioned version of the left hand side. The Richardson iteration for this preconditioned problem has the following form.

### v2

@@ 3.3 Schwarz Neural Inference
Normalization. During inference on an arbitrary decomposed subdomain, the range of geometry and boundary conditions may differ from that of the generated training data. We thus leverage the symmetry properties of PDEs to handle this mismatch. More specifically, we can apply problem-dependent transformations $T_{k}$ (e.g., spatial translation/scaling and value shift/scaling) to map a local input $(\Omega_{k},B_{k}^{n},H_{k})$ that lies outside the training range back into the training range. When the PDE symmetry acts on additional inputs (e.g., coefficient fields, source terms, parameters), the transformation is extended to $H_{k}$ accordingly. After neural operator inference, we map the predicted local solution back via an inverse transformation $\tilde{T}_{k}$. We implement these as preprocessing and postprocessing steps in the inference pipeline (see Appendix E).
Time Complexity. Suppose the single inference time of local operator and the number of iterations are denoted as $b$ and $N$. Let $v,e,K$ denote the number of vertices, edges and subdomains respectively. Our Algorithm 1 consists of two main parts: mesh partition using the METIS algorithm, the time complexity of which is approximately $O(v+e+K\log K)$ (Karypis and Kumar, 1997); and an additive-Schwarz-style iterative scheme with a time complexity roughly $O(bKN)$. Therefore, the overall time complexity of our algorithm can be approximated as $O(v+e+K\log K+bKN)$ [^1] With neural operators implementing linear transformers, e.g., GNOT applied in this work, $b=O(\frac{v}{K}).$. In practice, $K$ may not be independent of $v$ and $e$. More vertices can sometimes lead to more partitions required depending on the property of mesh. While providing an exact time complexity analysis for FEM can be challenging due to the complexity and variability of different problem setups, it is worth noting that FEM is generally considered to be computationally demanding.
@@ 3.4 Theoretical Results
Assume the operator $\mathcal{L}$ in Eq. 1 is self-adjoint and coercive elliptic (Mathew, 2008). Let $u^{n}$ be the iterates of the classical additive Schwarz-Richardson iteration in Eq. 2 and let $\tilde{u}^{n}$ be the iterates of SNI in Algorithm 1, initialized with the same $u^{0}=\tilde{u}^{0}$.
Let $\mathcal{S}_{k}$ denote the exact local solution operator on $\Omega_{k}$, and define the learned (pre/post-processed) local solver $\tilde{\mathcal{S}}_{k}\triangleq\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}$. Assume:
  • (Classical contractivity) the classical iteration mapping is contractive in a norm $\|\cdot\|$ with rate $0<\rho<1$;
  • (Uniform local error) $\|\tilde{\mathcal{S}}_{k}(\cdot)-\mathcal{S}_{k}(\cdot)\|\leq c$ for all $k$;
  • (Overlap factor) each degree of freedom is covered by at most $t$ subdomains.
Then we have:
  • Convergence: SNI converges to a fixed point $\tilde{u}^{*}$.
  • Error bound: for all $n$, $\|\tilde{u}^{n}-u^{n}\|\leq c^{\prime}$, and $\|\tilde{u}^{*}-u^{*}\|\leq c^{\prime}$, where one may take
    $c^{\prime}\;=\;\frac{\tau t}{1-\rho}\,c.$
The theorem suggests that if our learned local operator maintains a uniform error bound, the algorithm converges and exhibits a minimal approximation error. See Appendix C for a proof. This result relies on the assumption on operator $\mathcal{L}$. In general, such convergence is not guaranteed and we empirically validate the effectiveness of our framework for nonlinear differential equation through experiment.
@@ 4.1 Experimental Setup
Datasets. To demonstrate the scalability and superiority of our method, we construct several datasets on multiple PDEs. We also extend our framework to a time-dependent problem, heat conduction. To aggregate training sets, we generate random simple polygons bounded by the unit square $0.5,0.5]^{2}\subset\mathbb{R}^{2}$. Boundary/initial conditions and coefficient functions are piecewise linear functions determined by random values within $[0,1]$. For each of the following problems, we test on datasets based on three different domains A, B and C shown in Figure 2. Details of these datasets are given in Appendix H.1.
@@
  • NonlinearLaplace2d: A nonlinear Laplace equation in 2d with pure Dirichlet boundary condition on various shapes.
Baseline. Our baseline is a direct inference of the trained neural operator on domains shifted and scaled to $0.5,0.5]^{2}$ with boundary/initial conditions and coefficient functions adjusted accordingly.
Evaluation Protocol. The evaluation metric we utilize is the mean $l_{2}$ relative error. See Appendix H.2 for details.
@@ 4.2 Main Results and Analysis
Equation Domain GNOT(%) SNI(%)
[…]
NonlinearLaplace2d A $\pm2 2$$\pm0$.4
[…]
@@
The main results for all datasets are shown in Table 1. More details and hyperparameters are summarized in Appendix H.2 due to limited space. Based on these results, we have the following observations.
@@
Figure 3: Comparison between the $l_{2}$ relative errors from SNI (blue), GNOT direct inference (orange) on Laplace2d-Dirichlet upon three domains (A, B and C) with different numbers of training samples. The results of SNI and GNOT direct inference are presented based on 100 inferences with different boundary conditions. The best validation errors during training (red) are also provided as a reference.
@@
Time-dependent Problems. There is a natural way to extend our framework to time-dependent problems (Li and Cai, 2015) where a space-time decomposition is constructed by taking the product of a spatial decomposition and a temporal decomposition. We train a neural operator that can predict heat conduction on multiple time steps and the same SNI is applied during inference on this 3d problem. Our framework works well on this problem and reduce prediction error by 54.1%-74.2%. This demonstrates the potential of our framework to handle time-dependent problems. We refer to Appendix D for detailed implementation.
Data Efficiency. The exploration results on data efficiency of SNI are shown in Figure 3, implying the following observations: (1) At all abundances of data, the $l_{2}$ relative errors of SNI are significantly lower than those of GNOT direct inferences; (2) Errors of SNI are comparable to or even lower than validation errors at large data volumes. (3) SNI requires much smaller datasets to achieve comparable results to GNOT direct inference. Overall, these results demonstrate that SNI has substantial advantages in terms of data efficiency. Our proposed framework possesses remarkable ability to extract more insights from limited data and scale more effectively as data volumes increase. More supplementary results are provided in Appendix H.4.
@@ 4.3 Ablation Experiments
Validation(%) Domain A(%) Domain B(%) Domain C(%)
[…]
@@
Hyperparameter Exploration. The number of partitions ($K$), the depth of extension ($d$) and step size ($\tau$) are the key main hyperparameters that can affect the performance of SNI. Based on the results presented in Figure 4, the factors analyzed have no significant impact on the accuracy of our algorithm, but they do influence the convergence rate. Specifically, increasing the number of partitions leads to a smaller $l_{2}$ relative error but slower convergence. Once the partition number surpasses 20, the algorithm’s final performances become comparable. Regarding the depth of extension, it does not affect the performance on the tested domains. The convergence curves for depth of extension 1, 2, 4, and 8 are nearly identical. When it comes to $\tau$, a larger value results in faster convergence. However, it is important to note that there exists a maximum limit $1/K$ beyond which $\tau$ cannot be set.
Data Augmentation Exploration. To explore the effects of data augmentation, we compare the performances of models trained with different degrees of data augmentation for Laplace2d-Dirichlet demonstrated in Table 2. For models trained without data augmentation, the variation of performances on different domains is large, ranging from $2.8\%$ to $4.4\%$. Specifically, it reports a $4.4\pm 1.6\%$ $l_{2}$ relative error on domain A, while this error can be reduced to $1.9\pm 0.4\%$ with a rotation+[0.8, 1] scaling augmentation. While rotation can generally be beneficial, the effectiveness of scaling can sometimes be limited or even detrimental. Hence, it is important to apply data augmentation with caution and consider its suitability for different types of PDEs.
Choice of neural operator architecture. To explore the choice of neural operator architecture in our framework, we train a Geo-FNO (Li et al., 2023) on Laplace2d-Dirichlet and apply SNI for inference on domains A, B and C to get $l_{2}$ relative error of $9\pm 3\%$, $13\pm 1\%$ and $13\pm 3\%$. This result is comparable to that achieved by SNI with GNOT and demonstrates that our proposed framework works with various choices of neural operator architecture. However, an error gap does exists between SNI with GNOT and Geo-FNO due to variations in their generalizability. This is also reflected in their respective best validation errors, as detailed in Appendix  H.4. Supplementary results on Darcy2d are also provided there.
@@ 5 Related Work
Operator Learning. The idea of operator learning is first introduced in Lu et al. (2019). This work proposes a notable architecture called DeepONet, which employs a branch network for processing input functions and a trunk network for handling query points. Adopting the trunk-branch architecture and utilizing the attention mechanism,  Hao et al. (2023) develops GNOT to handle irregular mesh, multiple input functions, and different input data types. The high accuracy and versatility makes GNOT the benchmark in our work. In the other direction, Fourier neural operator (FNO) (Li et al., 2020b) leverages the Fast Fourier Transform (FFT) to learn operators in the spectral domain, and achieves a favorable trade-off between cost and accuracy. Variants of FNO are proposed to reduce computational cost (FFNO in Tran et al. (2021)), handle irregular mesh (Geo-FNO in Li et al. (2023)), and improve expressivity (UFNO in Wen et al. (2022)).
Methods to Deal with Complex Geometry. Several approaches have been proposed to tackle the challenge of complex geometry and save the process efforts in operator learning. One encoder-process-decoder framework called CORAL (Serrano et al., 2023) is able to encode a complex geometry into a lower dimensional representation to save the computational efforts and solve different types of problems. In (Wu et al., 2024), one mechanism called physics attention is proposed to aggregate complex input geometry and functions into several physics-aware tokens to reduce the number of tokens to deal with. AROMA (Serrano et al., 2024) introduces a diffusion refiner in latent space to solve temporal problems with complex geometries.
@@
Domain Decomposition Methods Applied in Deep Learning. In general, the integration of deep learning and DDMs can be categorized into two groups (Heinlein et al., 2021; Klawonn et al., 2024). The first category involves using deep learning techniques to improve the convergence properties or computational efficiency of DDMs. For instance, Mao et al. (2024) proposes to combine operator learning with DDMs on uniform grids in order to accelerate traditional DDMs. Several methods (Heinlein et al., 2020; 2019) have also been proposed to reduce the computational cost in adaptive FETI-DP solvers by incorporating deep neural networks while ensuring the robustness and convergence behavior. The second category is centered around the substitution of subdomain solvers in DDMs with neural networks. There have been multiple endeavors to employ PINNs (XPINNs (Jagtap and Karniadakis, 2020), parallel inference with cPINNs and XPINNs (Shukla et al., 2021)) or Deep Ritz methods as alternatives to subdomain solvers or discretization techniques in traditional DDMs (Li et al., 2020a; 2019; Jiao et al., 2021). These approaches leverage the universal approximation capabilities of neural networks to represent solutions of PDEs, subject to specific assumptions regarding the activation function and other factors.
@@
Data Augmentation Techniques in Operator Learning. Different types of data augmentations are proposed to improve the generalization capabilities in operator learning. A Lie point symmetry framework is introduced in Brandstetter et al. (2022), which quantitatively derives a comprehensive set of data transformations, to reduce the sample complexity. Motivated by this approach,  Mialon et al. (2023) learn general-purpose representations of PDEs from heterogeneous data by implementing joint embedding methods for self-supervised learning. An alternative research approach (Fanaskov et al., 2023) introduces a computationally efficient augmentation strategy that relies on general covariance and straightforward random coordinate transformations. In general, applying data augmentation techniques for PDE operator learning can be challenging due to the unique nature of PDE theory.
@@ 6 Conclusion and Future Works
Reproducibility statement
Detailed descriptions of the experimental setup, task definitions, and evaluation metrics are provided in section 4 and Appendix H. Source code is attached in the submission.
Appendix A Background on Domain Decomposition
@@
In this work, we decompose our domain into subdomains and adopt the hybrid formulation of Eq. 1 following Mathew (2008, Section 1.1). A decomposition of $\Omega$ is a collection of open subregions $\{\Omega_{k}\}_{k=1}^{K}$, $\Omega_{k}\subseteq\Omega$ for $k=1,\dots,K$ such that $\bigcup_{k=1}^{K}\overline{\Omega_{k}}=\overline{\Omega}$. This decomposition is referred to as non-overlapping if in addition, $\Omega_{i}\cap\Omega_{j}=\varnothing$ for any $i\neq j$. Alternatively, an overlapping decomposition is one satisfying $\bigcup_{k=1}^{K}\Omega_{k}=\Omega$. Typically, a non-overlapping decomposition is one where subdomains do not intersect with each other in the interior while an overlapping decomposition constructed in practice has overlapping neighboring subdomains.
@@
Given a decomposition of $\Omega$, a hybrid formulation of Eq. 1 is a coupled system of local PDEs on subdomains $\Omega_{k}$ equivalent to Eq. 1 satisfying two requirements. First, the restriction $u_{k}(x)$ of the solution $u(x)$ of Eq. 1 to each domain $\Omega_{k}$ must solve the local PDE, thus ensures that the hybrid formulation is consistent with the original problem in Eq. 1. Second, the hybrid formulation must be well posed as a coupled system of PDEs in the sense of Evans (2022), i.e. its solution must exist, be unique and depend continuously on given input function and boundary/initial conditions. Intuitively, a hybrid formulation consists of a local problem posed on each subdomain and matching conditions that couples the local problems.
@@
In this work we focus on the earliest and most elementary formulation termed Schwarz hybrid formulation (Mathew, 2008, Section 1.2) based on overlapping decomposition and is applicable to a wide class of self-adjoint and coercive elliptic equations. Given an overlapping decomposition, $\partial\Omega_{k}$ can be decomposed into two disjoint parts. One (possibly empty) part $\Gamma_{k}=\partial\Omega_{k}\cap\partial\Omega$ is located in the boundary of $\Omega$ and the global boundary condition should be imposed. The other part $B_{k}=\partial\Omega_{k}\cap\Omega$ is a nonempty artificial boundary from the overlapping decomposition and a Dirichlet boundary condition from the coupling of local problems is imposed.
@@
We refer to Mathew (2008) for a strict definition. As an illustrative example, assume we have an overlapping decomposition with $K=2$ and consider as the original problem Laplace equation with mixed Dirichlet and Neumann boundary conditions. The following coupled system of two local PDEs is a Schwarz hybrid formulation of the original problem and solving the original equation is equivalent to solving this coupled system.
[…]
@@
Based on the Schwarz hybrid formulation, there are various iterative schemes with different parallelism and convergence rate. In the subsequent discussion, our focus is primarily on introducing the additive Schwarz methods (ASM). The ASM is a highly parallel algorithm (Mathew, 2008) in solving the coupled system from Schwarz hybrid formulation. We briefly introduce ASM with finite element methods and refer to Gander and others (2008) and Mathew (2008) for details.
@@
Assume that under weak formulation of Eq. 1 and finite element space $V$, Eq. 1 has the form $Au=f$ where $A$ is the stiffness matrix. Given an overlapping decomposition $\{\Omega_{i}\}_{k=1}^{K}$ compatible with the finite element space on $\Omega$, we have $V=\sum_{k=1}^{K}V_{k}$ as sum of local finite element subspaces $V_{k}$ on $\Omega_{k}$ and we can define local stiffness matrices $A_{k}:V_{k}\to V_{k}$, restriction operators $\{R_{k}\}_{k=1}^{K}$ restricting $V$ to $V_{k}$ and extension operators $\{R_{k}^{\intercal}\}_{k=1}^{K}$ extending $V_{k}$ to $V$ by zeros extension. We then define operators $P_{k}:V\to V$ by $P_{k}=R_{k}^{\intercal}A_{k}^{-1}R_{k}A$. Additive Schwarz operator is then defined as the sum $P_{\text{ad}}=\sum_{k=1}^{K}P_{k}$. This operator can be show to be self-adjoint and coercive and we have the following equivalence.
@@
We note that the right hand side of Eq. 4 is a preconditioned version of the left hand side. The Richardson iteration for this preconditioned problem has the following form.